OpenStudy (anonymous):

Factor completely 1. a^3-ab+a^2-b 2. g^5-g 3. x^3-27y^9 4. 64x^6+8n^3 Factor completely 1. a^3-ab+a^2-b 2. g^5-g 3. x^3-27y^9 4. 64x^6+8n^3 @Mathematics

OpenStudy (turingtest):

1.(a^2-b)(a+1) 2.g(g^4-1) 3.(x-3y^3)(x^2+3xy^3+y^6) (difference of cubes) 4.(4x^2+2n)(16x^4-8x^2n+4n^2) (sum of cubes)

OpenStudy (anonymous):

g(g^4-1) - Can't you do the difference of 2 perfect squares? g(g^2-1)(g^2+1) ?

OpenStudy (turingtest):

yes, my bad I spaced out on that one :) you can actually do it twice: 3.g(g-1)(g+1)(g^2+1)

OpenStudy (anonymous):

And for this: (4x^2+2n)(16x^4-8x^2n+4n^2) when you put -8x^2n is that what you meant?

OpenStudy (turingtest):

yes, sum of cubes is a^3+b^3=(a+b)(a^2-ab+b^2) here you have a=4x^2 b=2n

OpenStudy (anonymous):

Wouldn't (a^2-ab+b^2) = (16x^4-(4x^2)(2n)+4n^2)?

OpenStudy (anonymous):

Not: (16x^4-8x^2n+4n^2)?

OpenStudy (turingtest):

yes, which simplifies to 16x^4-8x^2n+4n^2 (4x^2)(2n)=8x^2n you could put parentheses I guess if it looks confusing 16x^4-(8x^2)n+4n^2

OpenStudy (anonymous):

Oh I got it, I thought you raised it to the power of 2n, its confusing on the computer

OpenStudy (anonymous):

Thanks, do you think you're up for 1 more? 15x^2-43x-380

OpenStudy (turingtest):

15x^2-43x-380 (3x-20)(5x+19) i cheated for that one, it looked annoying. http://www.wolframalpha.com/input/?i=factor+15x^2-43x-380

OpenStudy (turingtest):

Wolfram can do most of those.

OpenStudy (anonymous):

Thanks so much, you're the best. I never heard of that and it will save me lots of time. :)